OpenStudy (anonymous):

Find all solutions in the interval [0,2pi]. tan x +sec x=1

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

help please!

OpenStudy (dumbcow):

\[\tan x + \sec x = 1\] \[\frac{\sin x}{\cos x} + \frac{1}{\cos x} = 1\] \[\frac{\sin x +1}{\cos x} = 1\] \[\sin x +1 = \cos x\] this only occurs when sin x = 0 or x=0

OpenStudy (anonymous):

could i ask you one more question please?

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

why does a calculator return an error when trying to find inverse sine of 1.055? is it because sine cant be more than 1?

OpenStudy (dumbcow):

correct range of sine and cosine is [-1,1]

OpenStudy (anonymous):

one more than im done. promise lol A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.02in./sec and that r = 36in. at time t = 0; Determine an equation that models the volume v of the balloon at time t, and find the volume v of the balloon at time t, and find the volume when t = 360 sec

OpenStudy (anonymous):

i got 4pi(36+0.02t)^3/3

OpenStudy (dumbcow):

you got it

OpenStudy (anonymous):

okay thank you!